In the Wall Bricks, Inc.’s balance sheet lists net fixed asset as $15 million. The fixed assets could currently be sold for $12 million. Wall Bricks’ current balance sheet shows current liabilities of $4 million and current assets of $10 million. If all the current accounts were liquidated today, the company would receive $5 million cash after paying the $2 million in liabilities. What is the book value of Wall Bricks’ assets today? What is the market value of these assets?
Answer #1

Book value of assets = Net fixed assets + current assets = 15 + 10 = 25 Million

Market value of assets = 12 + (5+2) = 19 Million